# Magnetic dipole-dipole transition for scintillation quenching

### v3

@@ I Introduction
The energy deposited by charge particles mainly causes the excitation and ionization of the solvent molecules. The energy is transferred from the solvent molecules to the solute molecules primarily through non-radiative transfer processes, for instance, the electric dipole-dipole transitions, which is also known as Förster resonance energy transfer, FRET [5], and the electric multipole-multipole resonance energy transfer suggested by Dexter [6]. The energy transfer may also occur through collision, for instance, the charge exchange effect [6]. The fluorescence of the solute molecules has a much longer attenuation length and detection efficiency for photomultiplier tubes.
However, when oxygen or molecules with heavy elements, for example, bromine, iodine, or lead, are dissolved in the solvent, a quenching effect always occurs [7]. Depending on the structure and property of the involved molecules, quenching can also happen through the non-radiative energy transfer and collision [5, 6] and the energy is dissipated to non-scintillating solutes. Quenching can also happen through intersystem crossing induced by spin-orbit coupling [8]. The intersystem crossing will cause the emission of phosphorescence instead of fluorescence.
In this work, a magnetic dipole-dipole transition mechanism is proposed as a scintillation quenching approach. The mechanism can also introduce an intersystem crossing and cause the loss of fluorescence. In section II, the energy transfer process in normal scintillation and the proposed quenching processes are described. Then the transition matrix is calculated in section III. The resonance condition and the influence of oxygen and heavy elements are presented in section IV. The result is discussed in section V, and the paper is summarized in the last section VII.
@@ II Scintillation and the proposed quenching effect
This paper is about the three component liquid scintillators, such as LAB+PPO with other research chemicals. The dominant and typical energy transfer process is plotted in Fig. 1, and the proposed new quenching approach is also shown in the figure. Other quenching mechanism is not the main topics of this paper.
In Fig. 1 and the paper, the solvent molecule is denoted as energy donor, $\mathrm{D}$, and the fluor solute molecule is denoted as $\mathrm{F}$. Donor’s electron is first excited from its ground state $^{1}\mathrm{D}_{0}$} to its first singlet excited state $^{1}\mathrm{D}^{*}$}. The prescript 1 refers to the singlet state, in which the spins of the outmost electron pair of the donor molecule are anti-parallel. The postscript 0 or * represents the ground or excited state, respectively.
@@
The de-excitation of $^{1}\mathrm{D}^{*}$} to $^{1}\mathrm{D}_{0}$} will cause $\mathrm{F}$ to be excited from its ground state, $^{1}\mathrm{F}_{0}$}, to first excited singlet state, $^{1}\mathrm{F}^{*}$}.
@@ III Magnetic dipole-dipole interaction
where $\mathbf{B}$ is the magnetic field caused by the donor electron and $\mathbf{m}_{\mathrm{A}}$ is the magnetic moment of the acceptor electron.
@@ IV Resonance condition
The energy difference, $\Delta U$, of an electron before and after spin flip in a magnetic field B is
@@
where $\Delta m_{z}$ is the change of magnetic moment in the direction of B, i.e., the $z$ direction, $q_{e}$ and $m_{e}$ are the electron’s charge and mass, respectively. Electron spin projection is $\pm 1/2$ in any direction, and the change between them is always 1. The result of $\Delta m_{z}$ is $({q_{e}}/{m_{e}})\hbar$ [13]. The magnetic field $B$ can be understood as the electron’s environmental magnetic field, contributed by other electrons. The gap between $^{1}\mathrm{D}^{*}$} and $^{3}\mathrm{D}^{*}$} is at the level of $1~\rm{eV}$ or smaller.
@@ IV.1 Oxygen
The oxygen, $\rm{O_{2}}$, is paramagnetic, i.e., attracted to a magnet, because there are two unpaired electrons in its outermost orbital. These electrons are ready to change their spin directions and stimulate the quenching.
@@ V Discussion
Motivated by FRET, the magnetic dipole-dipole interaction mechanism is proposed. In a fermion-antifermion system, magnetic dipole transition is quite feasible and known as M1-type decay [9, 16]. For example, these can be found in the decay of charmonium systems [17], mesons with a charm quark and an anti-charm quark, $J/\psi\to\gamma+\eta_{c}$, and $\psi(2S)\to\gamma+\eta_{c}^{\prime}$. The $\rm{J^{P}}$ (spin and parity) of $J/\psi$ and $\psi(2S)$ are both $1^{-}$, and that of $\eta_{c}$ and $\eta_{c}^{\prime}$ are both $0^{-}$. In view of particle physics, ${{}^{1}\mathrm{D}^{*}}\to{{}^{3}\mathrm{D}^{*}}$ processes occur naturally.
@@
From the point of view of quantum electrodynamics, one virtual M1-type photon is emitted from the donor electron and absorbed by the acceptor electron [9, 16, 18]. The parity of the M1-type photon is +1. The electric dipole radiation is mediated by a virtual E1-type photon, which has a parity of -1. Parity conservation requires there is no mixing between these two processes. The proposed magnetic dipole-dipole interaction is different from FRET.
@@ VI Possible experimental tests
One can think of using the known instrument, electron-paramagnetic resonance, EPR, to examine each interested liquid scintillator component and to search for free radicals, i.e., lone electrons. To see the free radical signals of LAB, the sample needs to be illuminated by ultraviolet light to excited states. One can also use the known organic with free radicals to check whether there is any enhanced quenching. But there are two difficulties in these tests. The first difficulty is to eliminate the interference of other quenching factors. The second difficulty is that the free radicals in the donor and acceptor may not satisfy the resonance condition, i.e., the magnetic fields of Eq. 10 for the lone electrons of the donor and acceptor do not match.
One can consider extending the known EPR to study the resonance condition. The current EPR devices apply two magnetic fields to the testing sample. One static magnetic field, $B_{0}$, is around 1 Tesla. The other one is an alternating magnetic field around tens giga Hz orthogonal to $B_{0}$. The current EPR instrument resonance signals occur at $B_{0}$, which completely covers the atom’s own magnetic field. The extension is to turn off the static magnetic field, i.e., with $B_{0}=0$, and to search for EPR resonance signals. The new measurement will tell us what the atom’s magnetic field in Eq. 10 is or what the range is.

### v5

@@ I Introduction
The energy deposited by charged particles mainly causes the excitation and ionization of the solvent molecules. The energy is transferred from the solvent molecules to the solute molecules primarily through non-radiative transfer processes, for instance, the electric dipole-dipole transitions, which is also known as Förster resonance energy transfer, FRET [11], and the electric multipole-multipole resonance energy transfer suggested by Dexter [8]. The energy transfer may also occur through collision, for instance, the charge exchange effect [8]. The fluorescence of the solute molecules has a much longer attenuation length and detection efficiency for photomultiplier tubes.
However, when oxygen or molecules with heavy elements, for example, bromine, iodine, or lead, are dissolved in the solvent, a quenching effect always occurs [6]. Depending on the structure and property of the involved molecules, quenching can happen through the non-radiative energy transfer and collision [11, 8], and the energy is dissipated to non-scintillating solutes. Quenching can also happen through intersystem crossing induced by spin-orbit coupling [20]. The intersystem crossing will cause the emission of phosphorescence instead of fluorescence.
In this work, a magnetic dipole-dipole transition mechanism is proposed as a scintillation quenching approach. The mechanism can also introduce an intersystem crossing and cause the loss of fluorescence. In Sec. II, the energy transfer process in normal scintillation and the proposed quenching processes are described. Then the transition matrix is calculated in Sec. III. The resonance condition and the influence of oxygen and heavy elements are presented in Sec. IV. The result is discussed in Sec. V, and a few experimental test proposals are raised in Sec. VI. The paper is summarized in the last section.
@@ II Scintillation and the proposed quenching effect
This paper is about the three-component liquid scintillators, such as LAB+PPO with other research chemicals. The dominant and typical energy transfer process is plotted in Fig. 1, and the proposed new quenching approach is also shown in the figure. Other quenching mechanism is not the main topics of this paper.
In Fig. 1 and the paper, the solvent molecule is denoted as energy donor, $\mathrm{D}$, and the fluor solute molecule is denoted as $\mathrm{F}$. The donor’s electron is first excited from its ground state $^{1}\mathrm{D}_{0}$} to its first singlet excited state $^{1}\mathrm{D}^{*}$}. The prescript 1 refers to the singlet state, in which the spins of the outmost electron pair of the donor molecule are anti-parallel. The postscript 0 or * represents the ground or excited state, respectively.
@@
The de-excitation of $^{1}\mathrm{D}^{*}$} to $^{1}\mathrm{D}_{0}$} will cause $\mathrm{F}$ to be excited from its ground state, $^{1}\mathrm{F}_{0}$}, to the first excited singlet state, $^{1}\mathrm{F}^{*}$}.
@@ III Magnetic dipole-dipole interaction
where $\mathbf{B}$ is the magnetic field caused by the donor electron, and $\mathbf{m}_{\mathrm{A}}$ is the magnetic moment of the acceptor electron.
@@ IV Resonance condition
The energy difference, $\Delta U$, of an electron before and after the spin flip in a magnetic field B is
@@
where $\Delta m_{z}$ is the change of magnetic moment in the direction of B, i.e., the $z$ direction, $q_{e}$ and $m_{e}$ are the electron’s charge and mass, respectively. Electron spin projection is $\pm 1/2$ in any direction, and the change between them is always 1. The result of $\Delta m_{z}$ is $({q_{e}}/{m_{e}})\hbar$ [9]. The magnetic field $B$ can be understood as the electron’s environmental magnetic field, contributed by other electrons. The gap between $^{1}\mathrm{D}^{*}$} and $^{3}\mathrm{D}^{*}$} for benzene is $0.96~\rm{eV}$ [17].
@@ IV.1 Oxygen
The oxygen, $\rm{O_{2}}$, is paramagnetic, i.e., attracted to a magnet, because there are two unpaired electrons in its outermost orbital. The ground state of $\rm{O_{2}}$ is a triplet [24]. These electrons are ready to change their spin directions and stimulate the quenching. The first excited state of $\rm{O_{2}}$ is a singlet, with the spin flipped for one electron. The energy level of the first excited state is 0.977 eV [24]. The mole fraction solubility of oxygen in water is $5\times 10^{-5} [22$] for the room temperature and the standard atmosphere pressure, and the mole fraction solubility of oxygen in benzene is $8.0\times 10^{-4}$ [23]. It can be seen that the energy gap is consistent with the energy gap between the benzene’s first excited singlet, $^{1}\mathrm{D}^{*}$}, and the first excited triplet, $^{3}\mathrm{D}^{*}$}, and, with the low solubility, a long-range interaction is expected.
@@ V Discussion
Motivated by FRET, the magnetic dipole-dipole interaction mechanism is proposed in this work. In a fermion-antifermion system, the magnetic dipole transition is quite feasible and known as M1-type decay [12, 10]. For example, these can be found in the decay of charmonium systems [3], mesons with a charm quark and an anti-charm quark, $J/\psi\to\gamma+\eta_{c}$, and $\psi(2S)\to\gamma+\eta_{c}^{\prime}$. The $\rm{J^{P}}$ (spin and parity) of $J/\psi$ and $\psi(2S)$ are both $1^{-}$, and that of $\eta_{c}$ and $\eta_{c}^{\prime}$ are both $0^{-}$. In view of particle physics, ${{}^{1}\mathrm{D}^{*}}\to{{}^{3}\mathrm{D}^{*}}$ processes occur naturally.
@@
From the point of view of quantum electrodynamics, one virtual M1-type photon is emitted from the donor electron and absorbed by the acceptor electron [12, 10, 13]. The parity of the M1-type photon is +1. The electric dipole radiation is mediated by a virtual E1-type photon, which has a parity of -1. Parity conservation requires that there is no mixing between these two processes. The proposed magnetic dipole-dipole interaction is different from FRET.
@@ VI Possible experimental tests
To avoid the quenching induced by the magnetic dipole-dipole transition, some study methods for oxygen gas [18] can be applied. It is important to first identify whether there is any absorption band around 1240 nm, i.e. 1 eV. Then one needs to identify whether the absorption band is from a spin-flip transition or a spin-conserving transition. If it is a spin-flip transition, it will be a target acceptor for the magnetic dipole-dipole transition.
Xenon is a good example, which has a high proton number, Z, and a moderate solubility in organic liquid scintillator, but a minor quenching effect. According to the reported information given by the KamLAND-zen experiment [1], 754 kg of xenon has been dissolved in liquid scintillator, and the mass fraction is 3.13%; no serious scintillation quenching has been reported. Checking the first few excited energy levels, one can find that the energy gap between the ground state and the first excited state has reached 8.3 eV [16]. It does not satisfy the resonance condition of electric or magnetic dipole-dipole transitions, so there is no serious quenching induced by electric or magnetic dipole-dipole transitions.
One can also think of using the known instrument, electron-paramagnetic resonance, EPR, to examine each interested liquid scintillator component and to search for free radicals, i.e., lone electrons. To see the free radical signals of LAB, the sample needs to be illuminated by ultraviolet light to excited states. One can also use the known organic with free radicals to check whether there is any enhanced quenching. But there are two difficulties in these tests. The first difficulty is to eliminate the interference of other quenching factors. The second difficulty is that the free radicals in the donor and acceptor may not satisfy the resonance condition, i.e., the magnetic fields of Eq. 10 for the lone electrons of the donor and acceptor do not match.
The known EPR can be extended to study the resonance condition. The current EPR devices apply two magnetic fields to the testing sample. One static magnetic field, $B_{0}$, is around 1 Tesla. The other one is an alternating magnetic field around tens giga Hz orthogonal to $B_{0}$. The current EPR instrument resonance signals occur at $B_{0}$, which completely covers the atom’s own magnetic field. The extension is to turn off the static magnetic field, i.e., with $B_{0}=0$, and to search for EPR resonance signals. The new measurement will tell us what the atom’s magnetic field in Eq. 10 is or what the range is.